OpenStudy (anonymous):

Medal will be given to the best answer. Which of the following is a solution to the equation y = 3x - 1 ? A. (4,1) B. (2,5) C. (4,3) D. (0, -3)

OpenStudy (anonymous):

B 3(2)-1=5

OpenStudy (anonymous):

Which equation matches the statement: ''The sum of -4x and 2 is 9''? A. -4x + 2 =9 B. -4x + 9 =2 C. -4x(2) = 9 D. -4x - 2 =9

OpenStudy (anonymous):

A. The sum of -4x and 2(-4x+2) is(=) 9 so -4x+2=9

OpenStudy (anonymous):

Thanks! Do you mind sticking around to help me with some more?

OpenStudy (anonymous):

yea sure

OpenStudy (anonymous):

x - 6 = -18 x = -24 x = -12 x = 12 x = 6

OpenStudy (anonymous):

ok so youre going to move the -6 over to the right side and you do that by adding 6 to the left side. so you should get x=-18+6 x=-12

OpenStudy (anonymous):

4x + 3 = 47 x = 11 x = 40 x = 44 x = 50

OpenStudy (anonymous):

4x+3=47 4x=47-3 4x=44 x=11

OpenStudy (anonymous):

1/3x = 12 x = 4 x = 15 x = 36 x = 39

OpenStudy (anonymous):

Sorry @JustinAA97 I was taking pond but I'm having computer problems and its taking longer than I excepted

OpenStudy (anonymous):

long*

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

1/3 of 12 is 4

OpenStudy (anonymous):

I have two more questions than I'll be out of your hair :)

OpenStudy (anonymous):

ok thats fine

OpenStudy (anonymous):

3x + 3 = 21 x = 5 x = 6 x = 18 x = 36

OpenStudy (anonymous):

and this one: 3/2x - 4 = 16 x = 20 x = 13 1/3 x = 15 1/3 x = 30

OpenStudy (anonymous):

3x+3=21 3x=24 x=8

OpenStudy (anonymous):

3/2x-4=16 3/2x=20 x=30

OpenStudy (anonymous):

That all?

OpenStudy (anonymous):

Yes, thank you so much! I'll give you a medal and I already became your fan but thanks again.

OpenStudy (anonymous):

Thank you